R. Bachimanchi, K. Baggett, M. Drury, A. Freyberger, R. Geng, C. Hovater, A. Hutton, L. King, G. Krafft, A. Kimber, R. Legg, C. Mounts, R. Nelson, T. Plawski, T. Powers, A. Seryi, and S. Suhring.
Agenda and Notes
1. Operations Status and Summary/Review (Freyberger)
Operations over the past two weeks have been relatively quite. The energy has been scaled back to 1000 GeV in NLINAC and 1020 GeV in SLINAC.
2. Cryomodule Shuffle Planning (Freyberg)
There have been several meetings where cryomodule shuffles for the summer have been planned out with the goal to get energy margin back, especially in the north linac. Attached are various presentations
3. Investigation of "Quench" Faults (Powers)
Many indicated quench faults are not real quenches because the field energy leaves the cavities too quickly. Some preliminary investigations
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?